OpenStudy (hang254):

Find the fourth roots of 256(cos 280° + i sin 280°).

OpenStudy (anonymous):

take the fourth root of 256, divide 280 by 4 that is the first one anyways

OpenStudy (anonymous):

use DeMoivre's theorem

OpenStudy (anonymous):

let x be the required root then x^4 = 256(cos 280° + i sin 280°) x^4 = 256(cos (2*k*pi+280° )+ i sin (2*k*pi+280° ))

OpenStudy (hang254):

doesn't the 4th root of 256 = 4

OpenStudy (anonymous):

so x = ( (256)^(1/4 )) (cos (2*k*pi+280° )+ i sin (2*k*pi+280° ))^(1/4) x = ( (256)^(1/4 )) (cos (2*k*pi+280° )/4+ i sin (2*k*pi+280° )/4) [k=0,1,2,3]

OpenStudy (hang254):

alright

OpenStudy (anonymous):

i hope u will be able to continue further...

OpenStudy (hang254):

thanks for the help, i'll refer back to here if i get stuck. Thanks!

OpenStudy (anonymous):

z = 256(cos280 + i sin280) = 256[cos(280+360k) + i sin(280+360k)] z^(1/4) = 4[cos(70+90k) + i sin(70+90k)} : k = 0, 1, 2, 3 z0 = 4[cos(70) + i sin(70)] z1 = 4[cos(160) + i sin(160)] z2 = 4[cos(250) + i sin(250)] z3 = 4[cos(340) + i sin(340)] No, cos(453) + i sin(453) is not an acceptable answer. 453 is coterminal with your result for k = 0. cos(453) = cos(93) and sin(453) = sin(93). All the angles in the polar representation should be reduced to the coterminal angle in the interval 0 ≤ Θ < 2π